1. Increasing the capacity just before a bottleneck operation will improve the output of the process.

True or false

2. Increasing the capacity of the bottleneck resource always increases the capacity of process.

True or false

3. If the process has an average inventory of 5 units and the flow time is 1 hr, than the actual cycle time is 12 minutes.

True or false
Answer #1

Answer 1 : False

Explanation : To improve the output of the process one must know and identify the value added activities, communicate changes effectively, map existing process etc.

Answer 2 : True

Explanation : When the capacity of the WIP in bottleneck is been increased the outcome and the capacity changes.

Answer 3 : True

Explanation : To calculate cycle time, use the formula : cycle time = 1 / throughput rate, where throughput rate = units produced/ time, hence : Throughput time = 5 / 60 min ( 1 hr = 60 minutes) = 0.0833

Now, cycle time = 1 / 0.08333 = 12 minutes.